# A Control Allocation Algorithm for Hypersonic Glide Vehicles with Input Limitations

### v2

@@ 5.1 Control Input-Based Weighting Matrices
The main motivation of the here proposed iterative control allocation approach is to consider soft constraints on the distribution of the control efforts. To do so, a meaningful design of the weighting matrices $\mathbf{W}_{m}(\bm{x},\bm{u})$ and $\mathbf{W}_{r}(\dot{\bm{u}})$ is leveraged. The magnitude weighting matrix $\mathbf{W}_{m}(\bm{x},\bm{u})$ is designed to allocate weights to control effectors based on their current measured deflection state $u_{i}$, as well as to use the weighting to prioritize control surfaces with less thermal loads acting on them. The overall matrix $\mathbf{W}_{m}(\bm{x}(t),\bm{u}(t))$ is defined as:
@@
This formulation assigns higher weights to effectors with larger deflection states, reflecting a strategic preference to distribute control efforts equally among available actuators. This approach has the operational benefit of reducing the tendency for each actuator to reach saturation.
To define $w_{m_{T},i}(\bm{x}(t))$ in a computationally efficient manner, we utilized insights from wind tunnel experiments conducted at the DLR Institute of Aerodynamics and Flow Technology [27]. Figure 10 shows an image of the wind tunnel test of the GHGV-2 at Mach 8.7, angle of attack $\alpha=10^{\circ}$, and sideslip angle $\beta=0^{\circ}$, with all flaps deflected by $20^{\circ}$.
Figure 10: Wind tunnel test of GHGV-2 with $M=8.7$, $\alpha=10^{\circ}$, and $\beta=0^{\circ}$; all flaps deflected by $20^{\circ}$ [27].
@@
where $w_{r,i}(\dot{u}_{i}(t-T))$ represents, for each control surface $i$, a rate-dependent weighting function defined as
$w_{r,i}=\frac{|\dot{u}_{i}(t-T)|}{\dot{u}_{\mathrm{crit},i}(\bm{x}(t),\bm{\sigma}(t))},$ (47)
@@
$\dot{u}_{\mathrm{crit},i}(\bm{x}(t))=\begin{cases}|\dot{u}_{\max,i}(\bm{x}(t),\bm{\sigma}(t))|,&\dot{u}_{i}(t-T)\geq 0,\\[6.0pt] |\dot{u}_{\min,i}(\bm{x}(t),\bm{\sigma}(t))|,&\dot{u}_{i}(t-T)<0.\end{cases}$ (49)
The carefully crafted weighting matrices $\mathbf{W}_{m}$ and $\mathbf{W}_{r}$ contribute to the proposed control allocation algorithm’s overall efficiency and stability, aligning with the discussed operational objectives. However, it is important to note that the proposed weighting functions $w_{m,i}(\bm{x}(t),u_{i}(t))$ and $w_{r,i}(\dot{u}_{i}(t-T))$ establish linear relationships with respect to the maximum and minimum values $u_{\max,i}(\bm{x}(t))$, $u_{\min,i}(\bm{x}(t))$, $\dot{u}_{\max,i}(\bm{x}(t))$, and $\dot{u}_{\min,i}(\bm{x}(t))$. Similarly, they assume an affine correlation between aerodynamic drag and thermal loads. While this linear approach ensures computational simplicity and efficiency, it inherently represents an idealized approximation. The proposed methodology simplifies the weighting function design but may not fully capture the nonlinear dependencies observed in practice. Future work could investigate incorporating nonlinear weighting schemes to better account for the interactions between deflection states, aerodynamic drag, and thermal effects while maintaining computational feasibility.
@@ 5.2 Designer Preference-Based Steady State Solution
The proposed IDCA approach allows it to consider a desired steady-state control input vector $\bm{u}_{s}(t)$. This section discusses how a good guess for $\bm{u}_{s}(t)$ is obtained based on an a priori known baseline control input $\bm{u}_{r}(t)$ and designer preferences in the control effector usage of the hypersonic glide vehicle. Even though the computed $\bm{u}_{s}(t)$ might not always provide a perfect solution to the requested virtual control command or in some cases might not be physically feasible under certain constraints, it serves as a desirable reference point or target for the redistributed control allocation algorithm. As discussed in Sect. 8, in a scenario where no external disturbances or model uncertainties are present, the baseline control input $\bm{u}_{r}(t)$ would be sufficient to maintain the HGV on its reference trajectory. However, given the inherent uncertainties and potential disturbances in hypersonic flight, the control system also requires an attitude feedback controller to provide additional corrections. This controller computes a virtual control adjustment $\Delta\bm{\nu}(t)$, which compensates for deviations from the nominal trajectory. Assuming an affine nature for the control input, the desired steady state control input can be computed as:
@@
Fig. 11 illustrates how the appropriate control surfaces are identified for an example of a virtual pitch command $\Delta\nu_{y}(t)$. For a positive pitch command $\Delta\nu_{y}(t)>0$, both the upper flaps $u_{1}$ and $u_{2}$ are identified as appropriate. Conversely, for a negative pitch command $\Delta\nu_{y}(t)<0$, $u_{3}$ and $u_{4}$ make the most sense from a flight dynamical perspective. This leads to the following sign-based conditionalization of $\mathbf{B}_{\nu_{y}}(\Delta\nu_{y}(t))$ in the pitch-related row:
@@
This approach yields a reasonable and operationally desired steady state, considering factors such as actuator use, drag reduction, and thermal load minimization. If $\bm{u}_{s}(t)$ is unfeasible or not a solution to $\bm{\nu}(t)=\mathbf{B}\bm{u}(t)$, the proposed redistributed approach will iterate to search for a viable solution close to $\bm{u}_{s}(t)$.
@@ 6 Simulation Results
A simulation-based analysis was carried out to examine the proposed IDCA algorithm. The flight dynamics model and the applied aerodynamic data set used for the time simulations are provided in a MATLAB/Simulink environment developed for the control design of hypersonic flight vehicles; see [8] for more details.
@@ 6.1 Comparison analysis on static moment commands
As discussed in Sect. 2.1, the presented PICA and RPICA algorithms are not necessarily capable of handling asymmetrical control input limitations. To illustrate this limitation, a simplified example of a constrained control allocation problem is considered and visualized in Fig. 12. The setup involves two control inputs $\mathbf{u}=[u_{1},\,u_{2}]^{T}$, where the admissible set is bounded by asymmetrical magnitude constraints $\mathbf{u}_{\min}=[0,\,0]^{T}$ and $\mathbf{u}_{\max}=[1.5,\,1.5]^{T}$. In this stationary example, rate constraints are neglected. The control effectiveness matrix is defined as $\mathbf{B}=[0.5,\,-0.5]$, and the demanded virtual control input is given by $\nu=0.5$.
@@ 7 Conclusion
In conclusion, this paper presented an iterative dynamic control allocation approach for HGVs, addressing key challenges such as asymmetric input limitations, actuator coupling, and thermal load management. Designed for computational efficiency, the approach enables real-time implementation, optimizing control distribution and reducing thermal stresses to prevent actuator saturation. Simulation results confirm its effectiveness in maintaining control authority across diverse mission profiles, meeting the stringent demands of HGV embedded systems.
Future work will incorporate state-dependent hard constraints, including limitations on achievable acceleration due to finite actuator forces. Additional studies will refine the cost function to better capture the nonlinear dynamics, further enhancing robustness and adaptability for hypersonic control applications.

### v3

@@ 5.1 Control Input-Based Weighting Matrices
The main motivation of the here proposed iterative control allocation approach is to consider soft constraints on the distribution of the control efforts. To do so, a meaningful design of the weighting matrices $\mathbf{W}_{m}$ and $\mathbf{W}_{r}$ is leveraged. The magnitude weighting matrix $\mathbf{W}_{m}$ is designed to allocate weights to control effectors based on their current measured deflection state $u_{i}(t)$, as well as to use the weighting to prioritize control surfaces with less thermal loads acting on them. The overall matrix $\mathbf{W}_{m}$ is defined as:
@@
This formulation assigns higher weights to effectors with larger normalized deflection states, thereby discouraging further usage of actuators that are already operating close to their admissible limits. It is worth noting that any standard $\ell_{2}$-based control allocation, such as the Moore–Penrose pseudoinverse, inherently promote an even distribution of control effort. This follows directly from their objective of minimizing the Euclidean norm of the control input vector, even in the absence of explicit weighting. However, those unweighted formulations do not account for the instantaneous actuator state or for state-dependent variations in admissible control authority. As a result, actuators that are already highly deflected or close to saturation may continue to be utilized disproportionately, potentially leading to premature magnitude or rate saturation. By contrast, the proposed weighting explicitly incorporates the current actuator deflection relative to its state- and operating-point-dependent maximum admissible value. This enables a dynamic redistribution of control effort toward effectors with greater remaining authority, rather than enforcing uniform usage in an absolute sense. The effect becomes particularly important under asymmetric or state-dependent actuator limitations, such as those induced by varying aerodynamic and thermal loads in hypersonic flight, where different control surfaces may experience substantially different admissible ranges. In this way, the weighting extends the equalizing tendency of standard $\ell_{2}$-based allocation by making it aware of actuator availability.
To address the thermal effects via $w_{m_{T},i}(\bm{x}(t))$ in a computationally efficient manner, we utilize insights from wind tunnel experiments conducted at the DLR Institute of Aerodynamics and Flow Technology [Hohn2022]. Figure 11 shows an image of the wind tunnel test of the GHGV-2 at Mach 8.7, angle of attack $\alpha=10^{\circ}$, and sideslip angle $\beta=0^{\circ}$, with all flaps deflected by $20^{\circ}$.
Figure 11: Wind tunnel test of GHGV-2 with $M=8.7$, $\alpha=10^{\circ}$, and $\beta=0^{\circ}$; all flaps deflected by $20^{\circ}$ [Hohn2022].
@@
where $w_{r,i}$ represents, for each control surface $i$, a rate-dependent weighting function defined as
$w_{r,i}=\frac{|\dot{u}_{i}(t)|}{\dot{u}_{\mathrm{crit},i}(\bm{x}(t),\bm{\sigma}(t))},$ (47)
@@
$\dot{u}_{\mathrm{crit},i}(\bm{x}(t))=\begin{cases}|\dot{u}_{\max,i}(\bm{x}(t),\bm{\sigma}(t))|,&\dot{u}_{i}(t)\geq 0,\\[6.0pt] |\dot{u}_{\min,i}(\bm{x}(t),\bm{\sigma}(t))|,&\dot{u}_{i}(t)<0.\end{cases}$ (49)
The carefully crafted weighting matrices $\mathbf{W}_{m}$ and $\mathbf{W}_{r}$ contribute to the proposed control allocation algorithm’s overall efficiency and stability, aligning with the discussed operational objectives. However, it is important to note that the proposed weighting functions $w_{m,i}$ and $w_{r,i}$ establish linear relationships with respect to the maximum and minimum values $u_{\max,i}(\bm{x}(t))$, $u_{\min,i}(\bm{x}(t))$, $\dot{u}_{\max,i}(\bm{x}(t))$, and $\dot{u}_{\min,i}(\bm{x}(t))$. Similarly, they assume an affine correlation between aerodynamic drag and thermal loads. While this linear approach ensures computational simplicity and efficiency, it inherently represents an idealized approximation. The proposed methodology simplifies the weighting function design but may not fully capture the nonlinear dependencies observed in practice.
@@ 5.2 Designer Preference-Based Steady State Solution
The proposed IDCA approach allows it to consider a desired steady-state control input vector $\bm{u}_{s}(t)$. This section discusses how a good guess for $\bm{u}_{s}(t)$ is obtained based on an a priori known baseline control input $\bm{u}_{r}(t)$ and designer preferences in the control effector usage of the hypersonic glide vehicle. Even though the computed $\bm{u}_{s}(t)$ might not always provide a perfect solution to the requested virtual control command or in some cases might not be physically feasible under certain constraints, it serves as a desirable reference point or target for the redistributed control allocation algorithm. As discussed in Section 4, in a scenario where no external disturbances or model uncertainties are present, the baseline control input $\bm{u}_{r}(t)$ would be sufficient to maintain the HGV on its reference trajectory. However, given the inherent uncertainties and potential disturbances in hypersonic flight, the control system also requires an attitude feedback controller to provide additional corrections. This controller computes a virtual control adjustment $\Delta\bm{\nu}(t)$, which compensates for deviations from the nominal trajectory. Assuming an affine nature for the control input, the desired steady state control input can be computed as:
@@
Figure 12 illustrates how the appropriate control surfaces are identified for an example of a virtual pitch command $\Delta\nu_{y}(t)$. For a positive pitch command $\Delta\nu_{y}(t)>0$, both the upper flaps $u_{1}$ and $u_{2}$ are identified as appropriate. Conversely, for a negative pitch command $\Delta\nu_{y}(t)<0$, $u_{3}$ and $u_{4}$ make the most sense from a flight dynamical perspective. This leads to the following sign-based conditionalization of $\mathbf{B}_{\nu_{y}}(\Delta\nu_{y}(t))$ in the pitch-related row:
@@
The resulting control input $\bm{u}_{s}(t)$ constitutes a reasonable and operationally meaningful steady-state reference. It is important to emphasize that $\bm{u}_{s}(t)$ is neither enforced as a hard constraint nor required to exactly satisfy $\bm{\nu}(t)=\mathbf{B}\bm{u}(t)$. Instead, it serves as a designer-preferred initial target for the subsequent redistributed dynamic control allocation, as shown by the closed-form solution in Eq. (36). The actual control input $\bm{u}(t)$ is obtained through an iterative redistribution process that explicitly accounts for feasibility, actuator limits, and the full control effectiveness matrix. If the reference $\bm{u}_{s}(t)$ is infeasible, the allocation naturally deviates from it and searches for a nearby admissible solution. From this perspective, the conditionalization can be interpreted as a structured initial guess that biases the solution toward desirable actuator usage patterns. Nevertheless, it must be noted that the induced sparsification implicitly introduces an allocation characteristic related to an $\ell_{1}$-type control allocation pattern. Such behavior is often advantageous in aerospace applications, as it promotes selective usage of the most effective control surfaces. However, as discussed in Section 2.1, purely $\ell_{1}$-based control allocation formulations tend to drive individual actuators toward their magnitude and rate limits, which can result in frequent constraint activation. In the proposed approach, this tendency is mitigated by embedding the designer-preferred steady-state solution as a soft reference within the proposed $\ell_{2}$-based dynamic control allocation approach. In this way, the sparsity-inducing properties of the steady-state preference are retained without being enforced as hard constraints, while the $\ell_{2}$-based dynamic control allocation law preserves compatibility with the full, non-simplified control effectiveness model as well as actuator magnitude and rate limitations.
@@ 6 Simulation Results
A simulation-based analysis was carried out to examine the proposed IDCA algorithm. The flight dynamics model and the applied aerodynamic data set used for the time simulations are provided in a MATLAB/Simulink environment developed for the control design of hypersonic flight vehicles; see [Autenrieb2021] for more details.
@@ 6.1 Comparison analysis on static moment commands
As discussed in Section 2.2, the presented PICA and RPICA algorithms are not necessarily capable of handling asymmetrical control input limitations. To illustrate this limitation, a simplified example of a constrained control allocation problem is considered and visualized in Fig. 13. The setup involves two control inputs $\mathbf{u}=[u_{1},\,u_{2}]^{T}$, where the admissible set is bounded by asymmetrical magnitude constraints $\mathbf{u}_{\min}=[0,\,0]^{T}$ and $\mathbf{u}_{\max}=[1.5,\,1.5]^{T}$. In this stationary example, rate constraints are neglected. The control effectiveness matrix is defined as $\mathbf{B}=[0.5,\,-0.5]$, and the demanded virtual control input is given by $\nu=0.5$.
@@ 7 Conclusion
In conclusion, this paper presented an iterative dynamic control allocation approach tailored to hypersonic glide vehicles and evaluated it in the context of asymmetric input limitations, actuator coupling, and state-dependent operating conditions. In contrast to pseudoinverse-based and redistributed pseudoinverse methods, the proposed formulation explicitly accounts for asymmetric magnitude and rate constraints and avoids degenerate or infeasible solutions under strong actuator coupling. Compared to quadratic programming–based control allocation approaches, the method significantly reduces computational complexity while retaining the ability to shape control distribution through soft constraints, making it suitable for real-time implementation in inner-loop flight control systems.
A key aspect of the proposed approach lies in the formulation of the control allocation problem itself. By appropriately designing state- and actuator-dependent weighting functions, the allocation process accounts for the instantaneous magnitude and rate states of the actuators as well as their directional control effectiveness. This enables a more balanced distribution of control effort and reduces the tendency of individual actuators to saturate, thereby preserving control authority under constrained conditions. In addition, the weighting function design provides a systematic mechanism to address thermal load effects. The use of drag-sensitive weighting functions allows thermal considerations to be incorporated implicitly into the allocation process without introducing additional hard constraints. By influencing control distribution based on drag generation characteristics, thermal loads and associated energy losses can be reduced while maintaining feasibility and compatibility with an iterative, approximate solution strategy.
Simulation results using the GHGV-2 model demonstrate that the proposed approach reliably respects actuator magnitude and rate limits, maintains control authority under realistic flight conditions, and offers a practical and computationally efficient alternative to existing state-of-the-art control allocation techniques for hypersonic applications.